OpenStudy (anonymous):

Suppose you are framing equilateral triangular garden beds. The side of each triangular bed measures 1.5 feet. How many complete triangular frames can you cut from a 200-inch piece of wood? (Hint: Remember to change measurements to the same units.)

OpenStudy (anonymous):

I'm not sure maybe @poopsiedoodle can help you

undeadknight26 (undeadknight26):

first convert the 200 inches to feet. i foot is 12 inches.

OpenStudy (anonymous):

soooooo , 17 ft..?

undeadknight26 (undeadknight26):

so: 200/12 = 16.66 now i think 1.5 x 3 = 4.5 16.66/4.5 = 3.70 which i think can be rounded to 4.@poopsiedoodle is this correct?

undeadknight26 (undeadknight26):

@poopsiedoodle is this right?

OpenStudy (mathstudent55):

I think it's better to work with inches.

OpenStudy (poopsiedoodle):

I'm not sure, but mathstudent seems to be on the job :p

OpenStudy (mathstudent55):

Each side of the triangle measures 1.5 ft. 3 sides measure 3 * 1.5 ft = 4.5 ft Now convert 4.5 ft to inches 4.5 ft * (12 in.)/(1 ft) = 54 in. Each bed will take 54 inches of wood. 200 inches divided by 54 inches will give you the number of frames you can do. 200/54 = 3.7037... Since the answer is 3.7..., that means you can frame three complete beds and seven tenths of the fourth bed. You don't have enough wood to fully frame the fourth triangular bed. Since the question is how many complete frames can you make, the answer is 3. This is one of the cases that you must round off down instead of to the nearest integer.

OpenStudy (anonymous):

thank you guys!c:

OpenStudy (mathstudent55):

wlcm